Changes between Version 72 and Version 73 of Status


Ignore:
Timestamp:
Jun 7, 2012 2:58:21 PM (6 years ago)
Author:
gaew
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • Status

    v72 v73  
    44
    55Please contact Dr Gary Llewellyn (Science / operations coordinator) if you have any questions or concerns: gaew@nerc.ac.uk (01452 859945 or 07919 697851)
     6
     7== ARSF Sitrep 07.06.11 ==
     8''During the rest of June the ARSF intend to collect data from the North of London (RG12-10), a location near Madrid, Spain (CR12-01), Harnhill & Diddington (GB12-06)  as well as several UK based projects that we were hopeing to support in May (GB09-04, GB12-10, BGS08-01 & BGS08-03/04/12-01). Of these RG12-10 holds the highest priority, though as always we are dependent on the weather!
     9''
     10
     11== ARSF Sitrep 23.05.11 ==
     12''Late on Sunday (20.05.12), Carl (Chief Pilot and Acting Head of the Facility) had to urgently return to the UK on a commercial flight. This related to the health of a member of his family, we wish him the very best of luck. However, James Johnson (our instrument operator) is also a commercial pilot which allowed our transit with the Dornier to continue.'' 
     13
     14''On Monday (21.05.12) we arrived in Athens and delivered the data collected from Milos and Santorini (and our data download station) to the Hellenic Military Mapping Service. Despite many repeated descriptions of our data media (Solid State Media in a Linex environment) it was quickly evident that none of the methods regarded as acceptable by the Greek Military would erase the data to the degree required by Greek Law. They regarded this as 'our' problem and were insistent on supervising any investigations into alternative methods. Significant assistance from Dr Grant at the ARSF-DAN yielded two possible solutions but these needed to be approved by senior Military Official before being used.''
     15
     16''On Tuesday morning we received news that one of the suggested methodologies would be accepted; we successfully used it on five SSDs and provided confirmation to the Military that these disks were now fully sanitised / erased of all data. On Tuesday the staff of the Hellenic Military Mapping Service were helpful, polite and willing to provide as must support as possible to facilitate data sanitation process. They advised that the sanitation / data removal from floating media disks in the past had taken days. Our process on our high data rate SSD disks took ~10 minutes / disk. I also had to opportunity to discuss the data processing that will have to be conducted at the Hellenic Military Mapping Service offices before these data will be released. It now seems possible that once they are satisfied that none of the frames and flight lines show any sensitive information that the raw data will also be released. ''
     17
     18''The two crew at the Hellenic Military Mapping Service offices met with the pilots (who organised the hotel check out) at the airport for a grueling four hour flight to Ljubljana (Slovinia), part way on the journey back to the UK. We will be back in the UK on Wednesday.''   
     19 
     20
     21== ARSF Sitrep 21.06.11 ==
     22''After two days (Friday and Saturday) of overcast conditions and rain Sunday was sunny and clear. The remaining flight lines over Milos were completed in the morning and the remaining Santorini site was completed in the afternoon. Indications are that all the data (VNIR, SWIR, LIDAR and digital photography) were good.''
     23
     24''On Monday we will set off early to Athens where all copies of the data will be transfered to the Greek Military. This transfer is likely to take at least one day.''
     25
     26== ARSF Sitrep 18.06.11 ==
     27''Yesterday (17.05.12), 11 survey lines (including a cross-line) were flown over the island of Milos in support of EM10-02. This constituted a third of the survey. It was suspended due to a build up of cloud and the last line will be repeated when we resume the survey later in the campaign. 165.8GB of data were collected. These included VNIR, SWIR, LiDAR and digital photography. We arrived on site at 10:27 hrs (local time) and the illumination is such that for most of the survey the Eagle data (VNIR) have a bandwidth of 1.25 nm (the best we have ever achieved). ''
     28
     29''Both sites (Santorini and Milos) have cloud cover today but conditions are better than forecast, probably due to a heavy thunderstorm last night that may have carried the forecast poor weather through the area at an accelerated rate. The crew are resting today in preparation for the resumption of data collection on Saturday and Sunday.''
     30
     31''Apart from a few summer colds, the crew are well and the instruments have performed well. A particular success is the enhanced data transfer speed of the data download station following an upgrade performed in the winter.''
     32
    633
    734== ARSF Sitrep 27.09.11 ==